OpenStudy (anonymous):

The three sides of a triangle are 5 inches, 8 inches and 9 inches. What is the best description for this triangle. acute equiangular obtuse right

OpenStudy (raden):

if the square of the longest side is < sum of square of the other then it is acute lets check together 9^2 = 81 5^2 + 8^2 = 25 + 64 = 89 9^2 < 5^2 + 8^2 so, it is an acute triangle